A chemist has a 90% solution of sulfuric acid and a 70% solution of the same acid. How many gallons of each must be mixed in order to make 200 gallons of a 75% solution
Thank you!

Answer by stanbon(75887)

A chemist has a 90% solution of sulfuric acid and a 70% solution of the same acid. How many gallons of each must be mixed in order to make 200 gallons of a 75% solution
-----------------
90% solution DATA:
amount= x gals. ; amount of active ingredient= 0.90x gals
-------------------
70% solution DATA:
amount= 200-x gals ; amount of active ingredient= 0.70(200-x)=140-0.7 gals
--------------------------
Mixture DATA:
amount = 200 gals ; amount of acive ingredient=0.75(200)= 150 gals
------------------------
EQUATION:
active + active = active
0.9x + 140-0.7x=150
0.2x=10
x=50 gals (Amount of 90% solution)
200-x=150 gals. (amount of 70% solution)
----------------------------------
Cheers,
Stan H.